OpenStudy (anonymous):

Write the expression (2 - 3i) + (-3 + 3i) in standard form.

OpenStudy (binary3i):

Z=(2-3i) + (-3+3i) = 2-3 +(-3+3)i -1

OpenStudy (anonymous):

how do you write it in standard form?

OpenStudy (anonymous):

That is in standard form. Unless you want \[e ^{i*\pi}\]

OpenStudy (anonymous):

okay, thanks.